OpenStudy (anonymous):

If each edge of a cube is double in lenght then the volume of the cube is multiplied by a factor of?

OpenStudy (anonymous):

8. For example, if you start each with 1 unit, and increase to 2, the volume is s^3 1*1*1=1 2*2*2=8 4*4*4=64